When 2.85 g2.85 g of a nonelectrolyte solute is dissolved in water to make 355 mL355 mL of solution at 28 °C,28 °C, the solution exerts an osmotic pressure of 937 torr.937 torr.

What is the molar concentration of the solution?

concentration:

MM

How many moles of solute are in the solution?

moles of solute:

molmol

What is the molar mass of the solute?

molar mass:

g/mol
Answer #1

Osmotic pressure = 937 torr = 1.23 atm

Temperature =T= 28oC =28+273 =301 K

Concentration of solution in molarity =C =n/V